# EgoCross: Benchmarking Multimodal Large Language Models for Cross-Domain Egocentric Video Question Answering

### v1

@@ 4 Experiments
4.1 Experimental Setup
In this section, we first outline the experimental setup in Section 4.1. The main results on our EgoCross benchmark are then presented in Section 4.2. In Section 4.3, we conduct a comparative analysis with in-domain benchmarks. Finally, Section 4.4 explores various methods to assess their potential and limitations in bridging the generalization gap.
Evaluated Models.
We select a diverse set of MLLMs spanning three categories to cover major technical paradigms: (1) To assess the current state-of-the-art performance, we include leading proprietary models: GPT-4.1 [16] and Gemini 2.5 Pro [17]. (2) For open-source general-purpose MLLMs, we consider Qwen2.5-VL (3B, 7B) [18], VideoLLaMA3 [20], and InternVL3 [19]. (3) To evaluate models tailored for egocentric understanding, we also include two egocentric-specialized models: EgoVLPv2 [9], and EgoGPT [10].
Evaluation Metrics.
Following prior works [3, 5, 7], we use standard accuracy metric for CloseQA, which is calculated as the percentage of correctly answered questions. For OpenQA, we employ a two-stage evaluation process: (1) a direct exact match between the generated and ground-truth answer, and (2) if no match is found, we adopt a LLM-as-a-Judge approach to evaluate semantic correctness.
Implementation Details.
All MLLMs are tested in a zero-shot setting with single-round inference. For video input, we extract frames at a fixed rate of 0.5 fps. For datasets that provide pre-sampled frames, we adhere to their original sampling frequency. No maximum frame limit is imposed to allow models to process the full temporal context. All experiments are conducted on NVIDIA A6000 GPUs.
4.2 Results on EgoCross
@@
Evaluation results are summarized in Table 2. We analyze the outcomes from four perspectives: (1) task-level challenges, (2) inter-domain variance, (3) model-wise performance, and (4) metric-type analysis.
Task-level Challenges. Most evaluated MLLMs struggle to perform well on our EgoCross benchmark, with average scores falling below 55% on CloseQA and below 35% on OpenQA. Considering that the random guess accuracy for CloseQA is 25%, these results suggest that the models indeed face substantial challenges in this benchmark.
Inter-Domain Variance. Across target domains, we observe varying levels of difficulty, ranging from relatively easy (Animal Perspective), middle-hard (Surgery) to particularly challenging (Extreme Sports, Industry). To further investigate inter-domain variance, we visualize t-SNE embeddings of EgoSchema and the four out-of-domain targets, using CLIP [43] as a modality-aligned feature extractor for both visual and textual representations (Figure 4).
@@
Figure 4: t-SNE visualization of text and visual features.
Model-wise Performance. The two proprietary MLLMs achieve the highest overall performance, with Gemini 2.5 Pro outperforming GPT-4.1. Following them are the open-source models, including Qwen2.5-VL, VideoLLaMA3, and InternVL3. Surprisingly, the egocentric-specific models (EgoVLPv2, and EgoGPT) perform the worst, despite being explicitly designed and trained on egocentric video data.
Metric-type Analysis. We further analyze the results under different evaluation metrics, namely CloseQA and OpenQA. Since CloseQA simplifies the task by providing explicit candidate answers, models naturally achieve higher accuracy in CloseQA compared to OpenQA. Additionally, we observe that CloseQA scores tend to be more stable across different MLLMs, while OpenQA is more sensitive to variations. For example, GPT-4.1 and Gemini 2.5 Pro achieve nearly identical scores on CloseQA (52.63 vs. 52.95), but differ noticeably on OpenQA (26.65 vs. 34.40).
@@ 4.3 More Analysis on Cross-Domain Gap
In Section 4.2, we demonstrate that domain gaps significantly contribute to the overall low performance. To further investigate this effect and highlight its unique presence in EgoCross, we compare model results on our benchmark and EgoSchema [5], a typical daily-life egocentric dataset featuring common activities like cooking and cleaning.
Figure 5: In-domain and cross-domain accuracy comparison. The results highlight the performance gap caused by domain shifts.
Results in Figure 5 reveal a consistent and significant performance drop across all comparable QA types. For instance, performance on action temporal localization drops from an impressive 92.31% on in-domain EgoSchema to just 34.13% on the novel domains of surgery, industry, and extreme sports in EgoCross. Similarly, next action prediction accuracy falls from 85.71% to 37.50%. The overall accuracy also drops from 73.58% to 43.14%, quantifying the substantial penalty incurred by the domain shift.
4.4 Pilot Studies
We proactively conduct several pilot studies to explore potential solutions for improving cross-domain egocentric QA. Specifically, we investigate three techniques: prompt learning, supervised fine-tuning (SFT), and reinforcement learning (RL). Since both SFT and RL require labeled data, we randomly split the initial test QA pairs into training and testing sets with a 70%:30% ratio. We adopt Qwen2.5-VL-7B as the baseline, and apply vLLM [44] for model acceleration. CloseQA results are shown in Table 3.
@@ 4.4 Pilot Studies
Table 3: CloseQA accuracy in pilot studies. ”+SFT” and ”+RL” denote supervised fine-tuning and reinforcement learning, respectively. * denotes the baseline without vLLM acceleration.
The results provide several insights: (1) Overall Trend. Each method, whether prompting (without labeled data) or SFT/RL (requiring labeled data), improves performance to some extent. (2) Impact of SFT. SFT boosts accuracy in domains like Industry (nearly 20% improvement). (3) Effectiveness of RL. RL shows the most significant improvement across all domains (an average increase of 22%).
@@ 5 Conclusion
In this work, we present EgoCross, a new benchmark for evaluating the cross-domain generalization ability of Multimodal Large Language Models (MLLMs) in egocentric video question answering. EgoCross comprises approximately 1k QA pairs based on video clips carefully collected and curated from four diverse and realistic domains: surgery, industry, extreme sports, and animal perspective. It spans a wide range of QA types and supports both CloseQA and OpenQA, enabling fine-grained evaluation. Beyond constructing the benchmark, we extensively evaluate current state-of-the-art MLLMs and find that, despite strong performance on existing benchmarks, they struggle to generalize to unfamiliar, domain-specific scenarios. Additionally, we further explore several potential techniques to improve cross-domain generalization. We believe that EgoCross, together with our experiments and analysis, offers a valuable foundation for future research on robust, domain-adaptive, and practically applicable egocentric video understanding.
Supplementary Material
6 Comparison with Existing Benchmarks
@@
7 More Implement Details
@@
7.1 Implement Details of Main Experiments
@@
7.2 Implement Details of Domain Gap Experiments
@@
7.2.1 Methodology for Semi-Automated Categorization
@@
7.2.2 Classification Rules and Keywords
@@
7.3 Implement Details Of Pilot Studies
@@
To ensure consistency and efficiency across experiments, we employ the vLLM framework [44] for both training and inference with the Qwen2.5-VL-7B model. While vLLM significantly accelerates these processes, we note a slight performance trade-off as shown in Table 3, which is an acceptable compromise, particularly for the extensive sampling required in RL training. All experiments were conducted on a server equipped with 8 NVIDIA H100 GPUs. The training setups for SFT and RL are inspired by the Video-R1 project, and we plan to release our code for reproducibility. For video inputs during training, we sample between 4 to 16 frames; sequences longer than 16 frames are truncated to 16. The inference settings for the pilot studies are kept consistent with those used for the main results.
@@
8 More Experiment Results
8.1 CloseQA and OpenQA evaluations
@@ C I L P C I L P I L P I L
9 More Details of Datasets

### v2

@@ 4 Experiments
4.1 Experimental Setup
Evaluated Models.
We select a diverse set of MLLMs spanning three categories to cover major technical paradigms: 1) To assess the current state-of-the-art performance, we include leading proprietary models: GPT-4.1 (Achiam et al. 2023) and Gemini 2.5 Pro (Comanici et al. 2025). 2) For open-source general-purpose MLLMs, we consider Qwen2.5-VL (3B, 7B) (Bai et al. 2025), VideoLLaMA3 (Zhang et al. 2025), and InternVL3 (Zhu et al. 2025). 3) To evaluate models tailored for egocentric understanding, we also include two egocentric-specialized models: EgoVLPv2 (Pramanick et al. 2023), and EgoGPT (Yang et al. 2025).
Evaluation Metrics.
Following prior works (Fan 2019; Mangalam et al. 2023; Plizzari et al. 2025), we use standard accuracy metric for CloseQA, which is calculated as the percentage of correctly answered questions. For OpenQA, we employ a two-stage evaluation process: 1) a direct exact match between the generated and ground-truth answer, and 2) if no match is found, we adopt a LLM-as-a-Judge approach to evaluate semantic correctness. Specifically, Qwen-MAX serves as the judge, providing a binary judgment (Correct/Incorrect) along with a detailed rationale for its decision. This ensures a robust assessment of semantic equivalence beyond simple string matching.
Implementation Details.
All MLLMs are tested in a zero-shot setting with single-round inference. For video input, we extract frames at a fixed rate of 0.5 fps. For datasets that provide pre-sampled frames, we adhere to their original sampling frequency (e.g., EgoSurgery at 0.5 fps and parts of CholecTrack20 at 1 fps). No maximum frame limit is imposed to allow models to process the full temporal context. All experiments are conducted on NVIDIA A6000 GPUs. More implementation details can be found in the Appendix.
4.2 Results on EgoCross
Evaluation results are summarized in Tab. 2. We analyze the outcomes from four perspectives: 1) task-level challenges, 2) inter-domain variance, 3) model-wise performance, and 4) metric-type analysis.
Task-level Challenges. Most evaluated MLLMs struggle to perform well on our EgoCross benchmark, with average scores falling below 55% on CloseQA and below 35% on OpenQA. Considering that the random guess accuracy for CloseQA is 25%, these results suggest that the models indeed face substantial challenges in this benchmark. Furthermore, excluding the top performance from proprietary models (Gemini 2.5 Pro and GPT-4.1), the remaining models perform notably worse, achieving less than 45% on CloseQA and under 20% on OpenQA. These observations collectively underscore both the difficulty and value of the proposed EgoCross benchmark, while also highlighting the current limitations of state-of-the-art MLLMs in handling cross-domain tasks.
Inter-Domain Variance. Across target domains, we observe varying levels of difficulty, ranging from relatively easy (Animal Perspective), middle-hard (Surgery) to particularly challenging (Extreme Sports, Industry). To further investigate inter-domain variance, we visualize t-SNE embeddings of EgoSchema and the four out-of-domain targets, using CLIP (Radford et al. 2021) as a modality-aligned feature extractor for both visual and textual representations (Fig. 4). The analysis highlights three key findings: 1) All target domains are clearly separated from EgoSchema and from each other in both visual and textual spaces, confirming the dataset’s cross-domain and diverse nature. 2) Within each of our target domains, textual features form sub-clusters, indicating the richness of our QA pairs. 3) The distributions help explain domain difficulty: Animal Perspective is closest to EgoSchema in both modalities, aligning with its relatively easier performance. In contrast, Industry and Extreme Sports are farthest, consistent with their higher difficulty. Surgery seems as a visual outlier but achieves relatively strong performance, suggesting that advanced MLLMs may be more robust to perceptual variation while still challenged by deeper semantic reasoning.
@@ 4.2 Results on EgoCross
Figure 4: t-SNE visualization of text and visual features. EgoCross domains are color-coded: Surgery (red), Industry (blue), ExtremeSports (green), Animal Perspective (orange) and Daily-activity (purple).
Model-wise Performance. As briefly discussed above, the two proprietary MLLMs achieve the highest overall performance, with Gemini 2.5 Pro outperforming GPT-4.1. This superiority is expected, given their access to large-scale training data and advanced algorithmic design. Following them are the open-source models, including Qwen2.5-VL, VideoLLaMA3, and InternVL3. Compared to proprietary models, they exhibit clear performance drops on both CloseQA and OpenQA tasks, highlighting the substantial gap that remains in tackling cross-domain egocentric understanding within the open-source community. Surprisingly, the egocentric-specific models (EgoVLPv2, and EgoGPT) perform the worst, despite being explicitly designed and trained on egocentric video data. Their failure, in contrast to general-purpose models, more clearly underscores the challenge of cross-domain generalization.
Metric-type Analysis. We further analyze the results under different evaluation metrics, namely CloseQA and OpenQA. Since CloseQA simplifies the task by providing explicit candidate answers, models naturally achieve higher accuracy in CloseQA compared to OpenQA. Additionally, we observe that CloseQA scores tend to be more stable across different MLLMs, while OpenQA is more sensitive to variations. For example, GPT-4.1 and Gemini 2.5 Pro achieve nearly identical scores on CloseQA (52.63 vs. 52.95), but differ noticeably on OpenQA (26.65 vs. 34.40). These observations confirm that OpenQA presents a more challenging setting, reflecting a common limitation of current MLLMs: their generative abilities are generally weaker than their judgment capabilities. By evaluating both OpenQA and CloseQA, we aim to comprehensively assess the generative and judgmental capacities of MLLMs in cross-domain scenarios.
In addition to dataset-level experiments and analysis, we conduct a more fine-grained evaluation across different QA types. Due to space limitations, detailed results and discussions are provided in the Appendix.
Figure 5: In-domain and cross-domain accuracy comparison across five QA types: Action Temporal Localization (ATL); Next Action Prediction (NAP); Interaction Identification (II); Special Action Identification (SAI); Action Sequence Identification (ASI). The results highlight the performance gap when evaluating on novel domains.
@@ 4.3 More Analysis on Cross-Domain Gap
In Sec. 4.2, we demonstrate that domain gaps significantly contribute to the overall low performance. To further investigate this effect and highlight its unique presence in EgoCross, we compare model performance between our benchmark and EgoSchema (Mangalam et al. 2023), a typical daily-life egocentric dataset featuring common activities like cooking and cleaning. To ensure a comparison across analogous QA types, we devise a semi-automated QA type categorization process to align EgoSchema’s QA pairs with our tasks. The detailed procedure is provided in the Appendix. Based on these aligned QA pairs, we evaluate Qwen2.5-VL, which achieves the best results among open-source MLLMs, on tasks from both EgoSchema (in-domain) and EgoCross (cross-domain) using the CloseQA protocol.
Results in Fig. 5 reveal a consistent and significant performance drop across all comparable QA types. For instance, performance on action temporal localization drops from an impressive 92.31% on in-domain EgoSchema to just 34.13% on the novel domains of surgery, industry, and extreme sports in EgoCross. Similarly, next action prediction accuracy falls from 85.71% to 37.50%. The overall accuracy also drops from 73.58% to 43.14%, quantifying the substantial penalty incurred by the domain shift. This task-level analysis reinforces our core finding: despite strong results on existing egocentric benchmarks, current MLLMs lack robustness when applied to unseen, domain-specific settings, a limitation largely overlooked in prior work.
@@
Table 3: CloseQA accuracy in pilot studies. “+SFT” and “+RL” denote supervised fine-tuning and reinforcement learning, respectively. * denotes the baseline without vLLM acceleration, it is marked in gray as vLLM acceleration causes slight performance degradation.
4.4 Pilot Studies
We proactively conduct several pilot studies to explore potential solutions for improving cross-domain egocentric QA. Specifically, we investigate three techniques: prompt learning, supervised fine-tuning (SFT), and reinforcement learning (RL). Since both SFT and RL require labeled data, we randomly split the initial test QA pairs into training and testing sets with a 70%:30% ratio. We adopt Qwen2.5-VL-7B as the baseline, and apply vLLM (Kwon et al. 2023) for model acceleration. CloseQA results are shown in Tab. 3.
The results provide several insights: 1) Overall Trend. Each method, whether prompting (without labeled data) or SFT/RL (requiring labeled data), improves performance to some extent. This suggests that refining prompt designs, expanding labeled data, and advancing algorithms are all promising directions for future exploration. 2) Impact of SFT. SFT boosts accuracy in domains like Industry (nearly 20% improvement). However, in some domains, such as Animal Per. no improvement is observed. This may stem from the inherently higher base performance in this domain, which is closer to natural domains. It could be caused by the limited number of training samples (Animal Per. has only 128 samples, far fewer than other domains.), which restricts the effectiveness of SFT. 3) Effectiveness of RL. RL shows the most significant improvement across all domains (an average increase of 22%). We attribute this to RL’s ability to learn from a broader range of interactions and feedback during training. The trial-and-error process enables the model to better handle longer sequences and more complex decision-making tasks, allowing it to dynamically adapt to the unique challenges of each domain.
@@ 5 Conclusion
In this work, we present EgoCross, a new benchmark for evaluating the cross-domain generalization ability of Multimodal Large Language Models (MLLMs) in egocentric video question answering. EgoCross comprises approximately 1k QA pairs based on video clips carefully collected and curated from four diverse and realistic domains: surgery, industry, extreme sports, and animal perspective, supporting both CloseQA and OpenQA for fine-grained evaluation. Our extensive evaluation reveals that current SOTA MLLMs struggle to generalize to these unfamiliar domains, despite strong performance on existing benchmarks. Additionally, we further explore several potential techniques to improve cross-domain generalization. We believe that EgoCross, together with our experiments and analysis, offers a valuable foundation for future research.
Acknowledgments
This work was supported by NSFC grant (No. 62136002 and 62477014), Ministry of Education Research Joint Fund Project (8091B042239), and Fundamental Research Funds for the Central Universities. Project supported by Shanghai Municipal Science and Technology Major Project (2025SHZDZX025G16).
Appendix A Comparison with Existing Benchmarks
@@
Appendix B More Implement Details
@@
B.1 Implement Details of Main Experiments
@@
B.2 Implement Details of Domain Gap Experiments
@@
Methodology for Semi-Automated Categorization
@@
Classification Rules and Keywords
@@
B.3 Implement Details Of Pilot Studies
@@
To ensure consistency and efficiency across experiments, we employ the vLLM framework (Kwon et al. 2023) for both training and inference with the Qwen2.5-VL-7B model. While vLLM significantly accelerates these processes, we note a slight performance trade-off as shown in Table 3, which is an acceptable compromise, particularly for the extensive sampling required in RL training. All experiments were conducted on a server equipped with 8 NVIDIA H100 GPUs. The training setups for SFT and RL are inspired by the Video-R1 project, and we plan to release our code for reproducibility. For video inputs during training, we sample between 4 to 16 frames; sequences longer than 16 frames are truncated to 16. The inference settings for the pilot studies are kept consistent with those used for the main results.
@@
Appendix C More Experiment Results
C.1 CloseQA and OpenQA evaluations
@@ C I L P C I L P I L P I L
Appendix D More Details of Datasets